# RG-Attn: Radian Glue Attention for Multi-modality Multi-agent Cooperative Perception

### v1

@@ IV-D Performance Analysis
Fusion Effectiveness: Both PTP and CoS-CoCo surpass the existing methods, as shown in Table. I, largely due to the effective cross-modality fusion capabilities of the RG-Attn module. PTP achieves dominant performance on both datasets, surpassing the previous SOTA method HEAL by 3.0% in DAIR-V2X AP30 and 1.9% in OPV2V AP70, extracting the maximum potential from the current RG-Attn design. This is attributed to the intra-agent fusion process, where camera semantic features are effectively “glued” to the intact sampled LiDAR-derived BEV feature map, fully transforming the camera features to enhance the BEV map. However, when conducting the cross-modality fusion in inter-agent stage of CoS-CoCo, the intersection area for sampling the BEV map to align with camera’s FOV might reach the boundary of LiDAR feature map, depending on the relative distance between two agents. Such cases can result in abnormal and fragmentary semantic information projection, leading to less effective fusion. The evident contrast can be observed in perception results visualization from Fig. 4. (b) to (d) consequently, with more objects identified and more accurate bounding boxes given incrementally.
@@
Ablation Component: The LiDAR-only cross-agent fusion setting for PTP or CoS-CoCo, representing the absence of RG-Attn for cross-modality fusion, leads to a performance degradation of 3.0% in AP30 and 2.7% in AP50 on DAIR-V2X, significantly higher than experimental error. Fig. 4. (a) can also reflect the essential functionality of RG-Attn, as the output BEV feature map exhibits a more distinct feature distribution.

### v2

@@ IV-D Performance Analysis
Fusion Effectiveness: Both PTP and CoS-CoCo surpass the existing methods, as shown in Table. I, largely due to the effective cross-modality fusion capabilities of the RG-Attn module. PTP achieves dominant performance on both datasets, surpassing the previous SOTA method HEAL by 3.0% in DAIR-V2X AP30 and 1.9% in OPV2V AP70, extracting the maximum potential from the current RG-Attn design. Such performance advantage of our proposed methods persists across varying numbers of participating agents, as shown in Fig. 4 and Fig. 5. This is attributed to the intra-agent fusion process, where camera semantic features are effectively “glued” to the intact sampled LiDAR-derived BEV feature map, fully transforming the camera features to enhance the BEV map. However, when conducting the cross-modality fusion in inter-agent stage of CoS-CoCo, the intersection area for sampling the BEV map to align with camera’s FOV might reach the boundary of LiDAR feature map or even no valid overlapping projection area, depending on the relative distance and direction of agents. Such cases can result in abnormal and fragmentary semantic information projection, leading to less effective fusion. The evident contrast can be observed in perception results visualization from Fig. 6 (b) to (d) consequently, with more objects identified and more accurate bounding boxes given incrementally.
@@
Ablation Component: The LiDAR-only cross-agent fusion setting for PTP or CoS-CoCo, representing the absence of RG-Attn for cross-modality fusion, leads to a performance degradation of 3.0% in AP30 and 2.7% in AP50 on DAIR-V2X, significantly higher than experimental error. Fig. 6 (a) can also reflect the essential functionality of RG-Attn, as the output BEV feature map exhibits a more distinct feature distribution.